OpenStudy (anonymous):

Simplify: log7^x + log7^8 - log7^2 =

OpenStudy (lgbasallote):

\[\huge \log_7 x + \log _7 8 - \log _7 2\] here's a hint \[\huge \log_a \left (\frac{bc}{d} \right) \implies \log_a b + \log_a c - \log_a d\] does that help?

OpenStudy (anonymous):

ok what would the next step be

OpenStudy (anonymous):

so it= log7=4x

OpenStudy (lgbasallote):

hmm?

OpenStudy (lgbasallote):

can you rewrite that?

OpenStudy (anonymous):

so you have log7 and divide the 8 by 2

OpenStudy (lgbasallote):

hmm...you sound right but i think it will be better if you draw what you mean so i can be sure if you're right or not

OpenStudy (anonymous):